OpenStudy (anonymous):

How many terms are there in the following sequence? 12,15,18,...,57

OpenStudy (anonymous):

16 terms. Increase each previous term by 3 until you reach 57, then count the total number of terms.

OpenStudy (anonymous):

OR use...

OpenStudy (tkhunny):

Insufficient information. You can ASSUME it is an arithmetic sequence, but there really is not enough information. Maybe it's 12, 15, 18, 32, 75, 318.4, 9, 57

OpenStudy (anonymous):

\[a _{n}=a _{1}+(n+1)d\]

OpenStudy (anonymous):

It's probably not insufficient information, though I get your point @tkhunny. Which math are you taking @worthb214!?

OpenStudy (tkhunny):

It IS insufficient information. The author of the problem should repent and do a better job at writing questions.

OpenStudy (anonymous):

Digressing back to the formula. Plug in the last term (57) for \[a _{n}\] and plug in the first term (12) for \[a _{1}\]. Then solve for 'n'.

OpenStudy (anonymous):

Its Math 102 for educators I don't know that is what the problem said.

OpenStudy (tkhunny):

For educators? That's the PERFECT forum for discussing the inadequacies of the problem statement. Excellent. Bring it up and show your strength.

OpenStudy (tkhunny):

It can be fixed with one word. How many terms are there in the following ARITHMETIC sequence? 12,15,18,...,57

OpenStudy (anonymous):

Like I said, it's 16. Use either of these two calculators to verify: http://www.basic-mathematics.com/arithmetic-sequence-calculator.html or http://www.calculator.net/number-sequence-calculator.html

OpenStudy (anonymous):

Did the question stem from a section titled "Arithmetic Sequences" or some derivation thereof?

OpenStudy (anonymous):

yeah that is what we are learning now is arithmetic and geometric sequences

OpenStudy (anonymous):

I mistyped the equation originally. It should read: \[a _{n}=a _{1}+(n-1)d\] 57 = 12 + (n - 1)3 Distribute the 3 and you'll have: 57 = 12 + 3n - 3 Combine like terms that are already on the same side of the equal sign: 57 = 9 + 3n Subtract 9 from both sides: 48 = 3n Divide both sides by 3 16 = n

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

This is a helpful page to checkout: http://www.regentsprep.org/regents/math/algtrig/atp2/arithseq.htm

OpenStudy (anonymous):

Take care!